# XENONnT WIMP Search: Signal & Background Modeling and Statistical Inference

### v1

@@ III.3 Neutron Background Model
The propagation of the neutrons in the XENONnT detector is simulated with the Geant4 toolkit [53, 54], where the recoil type and energy deposition per interaction in the target are recorded. We compute the number of photons and electrons for a given interaction via the custom-developed epix package [55], which utilizes the energy-dependent LXe response derived from the SR0 calibrations, shown in Fig. 3. This package also handles the clustering of the individual energy-depositing steps at the LXe microphysics scale before the quanta generation. The photons and electrons produced by epix are passed to the waveform simulator (WFSim) [56], which computes the S1 and S2 signals up to the waveform level, by means of a precise set of simulations- and data-driven corrections which characterize the XENONnT detector response. The event-by-event simulated waveforms share the same data structure as the science data after applying PMT and DAQ effects, which allows us to process them with the same software used for the real data (straxen [57]), as well as to apply the same data selections. The final SS neutron rate arises from weighting the rates obtained via the entire waveform simulation pipeline with the specific activities of the corresponding material and isotopic neutron yield.
@@
With this agreement between simulations and calibration data, we decided before unblinding the WIMP ROI to proceed with the sideband unblinding of the events tagged by the neutron veto. Initially, this confirmed the simulation-driven neutron background prediction. However, a mistake in the definition of the neutron veto time window was found after the unblinding of the WIMP ROI. After fixing this issue, a mismatch was found, with the neutron background rate being larger than predicted in the ROI [5]. We therefore decided to constrain the neutron background rate in a purely data-driven way based on the aforementioned sideband unblinding with the correct neutron veto tagging window. The results of the sideband unblinding are shown in Fig. 7. Based on the multiple-to-single scatter ratio of 2.2, the neutron veto tagging efficiency, and the three observed MS events and one SS neutron event tagged by the neutron veto in the fiducial volume, a data-driven prediction of events was derived for the SR0 exposure. We used the simulation-driven ratio between normal neutron background and neutron+X events to estimate the neutron+X event rate.
@@ III.5 Accidental Coincidence Background Model
Validation performed on calibration data provides the rate and shape uncertainty of the AC model. For this, we compared the predicted and observed AC events in a dedicated dataset by performing both 1D and 2D GOF tests with equiprobable bins in all relevant parameters. The AC prediction is provided by the AC modeling method discussed above, and the AC dataset is selected by inverting the data selections to avoid other physical events. Because of the large statistics of 37Ar calibration data, it delivers the most stringent constraint on the AC model. 37Ar events with S2 areas smaller than 400 PE are selected to test the AC model. The predicted number of AC events was 731.6, while 733 events were observed in the data. The statistics in the AC model is very large so the uncertainty of the predicted AC rate is neglected in the following GOF tests. The 1D GOF tests in S1, S2, , , and the 2D test in S1–S2 all yield p-values between 0.05 and 0.95. The result of the 2D test with a p-value of 0.61 is shown in LABEL:fig:ac_${}^{37}$Ar. The model was further validated with events removed by selections targeting the AC background inside the WIMP ROI. Similar tests were also performed on 220Rn calibration data. All these tests show the model and data to be compatible.
To suppress AC events, a selection criterion based on a gradient-boosted decision tree (GBDT) classifier utilizing the S2 pulse shape and the drift time was developed. Due to an insufficient model for the S2 pulse shape near the perpendicular wires, the GBDT classifier is only applied in the region at least 4.45 cm far from the wires (far-wire region), while in the region near the wires (near-wire region), a data-driven S2 shape selection is applied instead. The resulting AC background rate per ton-year in the near-wire region is about 5.6 times higher compared to the rate in the far-wire region. Consequently, the modeling of the TPC response is split into two parts. Due to an over-fitting issue found in the GBDT training process, we conservatively estimated the rate uncertainty to be 30% for both the near- and far-wire regions. The expectation value of AC events in the WIMP search dataset in the ROI is .
@@ III.6 Surface Background Model
In XENONnT, the TPC wall is made of PTFE, which is known to accumulate isotopes from the decay chain of atmospheric 222Rn, which decays down to 210Pb, an isotope with a half-life of 22.2 years. Radioactive decays on the wall surface can result in events with reduced S2 signals due to charge losses, which gives rise to a background that can leak into the WIMP signal region. Three components contribute to the surface background in the WIMP ROI: the two -decays of 210Pb and 210Bi, and the recoiling 206Pb following the -decay of 210Po. Due to the complexity of electric field conditions near the surface and the loss of ionization electrons to the detector walls, we employ a data-driven approach to model this background component in the space of cS1–cS2––.
The surface background model in the space of – was developed using 210Po -events. These events are mono-energetic (5.4 MeV) and thus easily identifiable through their characteristic S1 signal, yet they present a wide range of S2 sizes due to variable charge loss to the walls. The and distributions of 210Po -events were seen to match those of the lower-energy -events. The radial profiles were modeled using a Student’s generalized skew-t distribution. The radial distribution was fitted independently for different S2 sizes as shown in Fig. 9 to account for the S2-size dependent position resolution [10]. The profile was modeled using a linear function, to account for lower rates of surface events observed at the bottom of the TPC, likely due to the increased charged insensitivity near the walls at the bottom of the detector [7].
The cS1–cS2 distribution was modeled using a 2D Gaussian adaptive kernel density estimation (aKDE), built using events reconstructed outside the detector walls. The resulting model was then validated against the events reconstructed between the walls and the fiducial volume, in order to rule out any dependence between cS1, cS2, and R. Figure 4 (lower left) shows the projection of the four-dimensional model on cS1–cS2 in the ROI. The absolute rate of events in the blinded region was inferred from the radial distributions of adjacent, non-blinded regions in cS1–cS2.
@@ IV.2 Nuisance Parameters and Constraints
| Signal efficiency | NR model uncertainty | ||
@@
In addition to our parameter of interest , we parameterize the space of background hypotheses using nuisance parameters in the likelihood function of Eq. 5. These parameters control both background rates and shape as well as the signal efficiency. In total, twelve nuisance parameters are defined, which are listed in Tab. 1. The uncertainties of the nominal expectation values correspond to the width of a Gaussian constraint term. The ancillary likelihood function is the product of all constraint terms. Most parameters are constrained via ancillary measurements, which are obtained from sidebands (e.g. data outside the ROI) or external measurements in combination with simulations. More details on the constraints on neutron, CENS, AC, and surface background rates were discussed in Sec. III. The ER rate is a free parameter in the fit and is fully constrained by the WIMP search data. The ER shape parameters obtained with PCA are constrained via the simultaneous fit of the calibration dataset via the term in Eq. 5.
@@ IV.3 Confidence Intervals and Discovery Significance
Knowing the distribution of under different hypotheses is essential for calculating discovery significances and confidence intervals. Due to the low rate of background resembling the signal, asymptotic formulae as discussed in [64] are not applicable. Instead, toy-MCs are used to estimate the distribution of the test statistic given that is the true cross section. We use the custom-developed blueice [65] framework to define the likelihood function. This Python package provides an efficient interpolation (“template morphing”) between PDFs evaluated for different discrete values of shape nuisance parameters. This allows shape parameters to be considered for which we have no analytical description.
For each signal model considered, we compute the discovery significance and the confidence interval by comparing the measured test statistic with the distribution under each hypothesis [66, 67]. Testing yields the (local) discovery significance, while confidence intervals are computed by finding the region where is rejected at a confidence level (CL) given the data, which is illustrated for three background-only toy-MCs (three black parabola-like lines) in Fig. 10 (top). The minimum of each curve corresponds to the respective best-fit value . Upper limits (ULs) and lower limits are obtained by finding the cross sections at which reaches the critical region. The threshold of the 90 CL critical region is precomputed as the 90th percentile of the test statistic distribution from toy-MCs with injected signal corresponding to the respective cross section. For low cross sections, it deviates from the asymptotic two-sided threshold indicated as a horizontal gray dotted line.
@@
Statistical fluctuations as well as mismodeling, such as overestimated background rates, can yield arbitrarily low ULs, which may result in the spurious exclusion of models beyond the experiment’s sensitivity. To mitigate this issue, various methods have been proposed [68, 69, 70]. The XENONnT SR0 WIMP search follows the recommendation of [46] to use power-constrained limits (PCL) [70]. In this method, a signal size threshold is selected, at which the experiment has a “rejection power” – the probability of excluding a given signal under the background-only hypothesis. If an UL falls below this threshold, the threshold value is reported instead. These thresholds correspond to the quantiles of the UL distribution used to compute the sensitivity band illustrated in Fig. 10. For instance, choosing sets the threshold to the -1 sigma quantile of the band, while truncates ULs at the band’s median. In [46], the fiducial choice of was suggested. However, the choice was erroneously based on the discovery power, which corresponds to the probability of rejecting the background-only hypothesis given an alternative hypothesis with a specific signal size. In the absence of an updated recommendation, we chose the very conservative choice of for the first WIMP search of XENONnT.
@@ IV.4 Goodness of Fit and Blinding Procedure
To verify that the best-fit models describe our WIMP search data well, we defined a GOF test before unblinding, and after studies to optimize the power to reject mismodeling. The test uses a binned Poisson likelihood with 15 equiprobable bins in the cS1–cS2 space, defined from the model being tested. To compute a p-value, the distribution of the test statistic under the best-fit hypothesis is derived through toy-MCs. Specifying the test and acceptance threshold (90% CL) before unblinding ensured that the results were not influenced by statistical fluctuations expected from the low-statistic WIMP search dataset. For the background-only fit we found a p-value of and for a fit with an additional WIMP signal the p-value is . Both indicate no strong mismodeling in the predefined parameter space. Using GOF tests with well-studied power to discover signal-like mismodeling was also designed to replace the “safeguard” ER shape parameter defined in [71], due to its computational cost and susceptibility to some kinds of mismodeling.
When searching for new signals in data, it is crucial to avoid the experimenter’s bias on the result [72]. In our WIMP search, we adopt a common strategy for bias mitigation: blinding the signal region until all steps of the data analysis are finalized. Initially, events in the ER and NR bands with a reconstructed ER energy below were blinded. This involved defining the ER and NR bands in cS1 and cS2 based on quantiles in cS2 for slices in cS1 of calibration data. In the first step, all events above and those above the sigma quantile of the ER band were unblinded for the analysis presented in [63]. The signal region for the WIMP search (indicated in Fig. 5) remained blinded until the analysis procedure was finalized. The unblinded SR0 WIMP search data showed no significant excess for any of the tested WIMP masses with local discovery p-values . For this reason, we reported new ULs on the spin-independent WIMP-nucleon cross section across WIMP masses ranging from to , with a minimum of at .
@@ Acknowledgements.
We gratefully acknowledge support from the National Science Foundation, Swiss National Science Foundation, German Ministry for Education and Research, Max Planck Gesellschaft, Deutsche Forschungsgemeinschaft, Helmholtz Association, Dutch Research Council (NWO), Fundacao para a Ciencia e Tecnologia, Weizmann Institute of Science, Binational Science Foundation, Région des Pays de la Loire, Knut and Alice Wallenberg Foundation, Kavli Foundation, JSPS Kakenhi and JST FOREST Program ERAN in Japan, Tsinghua University Initiative Scientific Research Program, DIM-ACAV+ Région Ile-de-France, and Istituto Nazionale di Fisica Nucleare. This project has received funding/support from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN. We gratefully acknowledge support for providing computing and data-processing resources of the Open Science Pool and the European Grid Initiative, in the following computing centers: the CNRS/IN2P3 (Lyon - France), the Dutch national e-infrastructure with the support of SURF Cooperative, the Nikhef Data-Processing Facility (Amsterdam - Netherlands), the INFN-CNAF (Bologna - Italy), the San Diego Supercomputer Center (San Diego - USA) and the Enrico Fermi Institute (Chicago - USA). We acknowledge the support of the Research Computing Center (RCC) at The University of Chicago for providing computing resources for data analysis. We are grateful to Laboratori Nazionali del Gran Sasso for hosting and supporting the XENON project.References
@@
- E. Aprile et al. [2024a] E. Aprile et al. (XENON collaboration), The XENONnT Dark Matter Experiment, preprint (2024a), arXiv:2402.10446 [physics.ins-det] .
@@
- E. Aprile et al. [2024b] E. Aprile et al. (XENON collaboration), XENONnT Analysis: Signal Reconstruction, Calibration and Event Selection; In preparation (2024b).
@@
- E. Aprile et al. [2024c] E. Aprile et al. (XENON collaboration), Erratum to XENON1T Dark Matter Data Analysis: Signal & Background Models, and Statistical Inference, Phys. Rev. D (2024c), in preparation.
@@
- Lindhard et al. [1963] J. Lindhard, M. Scharff, and H. E. Schioett, Range concepts and heavy ion ranges (notes on atomic collisions, II), Kgl. Danske Videnskab. Selskab. Mat. Fys. Medd. 33, (1963).

### v2

@@ III.3 Neutron Background Model
The propagation of the neutrons in the XENONnT detector is simulated with the Geant4 toolkit [55, 56], where the recoil type and energy deposition per interaction in the target are recorded. We compute the number of photons and electrons for a given interaction via the custom-developed epix package [57], which utilizes the energy-dependent LXe response derived from the SR0 calibrations, shown in Fig. 3. This package also handles the clustering of the individual energy-depositing steps at the LXe microphysics scale before the quanta generation. The photons and electrons produced by epix are passed to the waveform simulator (WFSim) [58], which computes the S1 and S2 signals up to the waveform level, by means of a precise set of simulations- and data-driven corrections which characterize the XENONnT detector response. The event-by-event simulated waveforms share the same data structure as the science data after applying PMT and Data Acquisition (DAQ) effects, which allows us to process them with the same software used for the real data (straxen [59]), as well as to apply the same data selections. The final SS neutron rate arises from weighting the rates obtained via the entire waveform simulation pipeline with the specific activities of the corresponding material and isotopic neutron yield.
@@
With this agreement between simulations and calibration data, we decided before unblinding the WIMP ROI to proceed with the sideband unblinding of the events tagged by the neutron veto. Initially, this confirmed the simulation-driven neutron background prediction. However, a mistake in the definition of the neutron veto time window was found after the unblinding of the WIMP ROI. After fixing this issue, a mismatch was found, with the neutron background rate being larger than predicted in the ROI [5]. We therefore decided to constrain the neutron background rate in a purely data-driven way based on the aforementioned sideband unblinding with the correct neutron veto tagging window. The results of the sideband unblinding are shown in Fig. 7. Based on the multiple-to-single scatter ratio of 2.2, the neutron veto tagging efficiency, and the three observed MS events and one SS neutron event tagged by the neutron veto in the fiducial volume, a data-driven prediction of events was derived for the SR0 exposure. We used the simulation-driven ratio between normal neutron background and neutron+X events to estimate the neutron+X event rate shown in Tab. 1.
@@ III.5 Accidental Coincidence Background Model
To suppress AC events, a selection criterion based on a gradient-boosted decision tree (GBDT) classifier utilizing the S2 pulse shape and the drift time was developed. Due to an insufficient model for the S2 pulse shape near the perpendicular wires, the GBDT classifier is only applied in the region at least 4.45 cm far from the wires (far-wire region), while in the region near the wires (near-wire region), a data-driven S2 shape selection is applied instead. The resulting AC background rate per ton-year in the near-wire region is about 5.6 times higher compared to the rate in the far-wire region. Consequently, the modeling of the TPC response is split into two parts.
Validation performed on calibration data provides the rate and shape uncertainty of the AC model. For this, we used dedicated AC-rich datasets, such as sidebands defined by inverting anti-AC selections. We compared the predicted and observed AC events by performing both 1D and 2D GOF tests with equiprobable bins in all relevant parameters. The AC predictions are provided by the AC modeling method discussed above. Because of the large statistics of 37Ar calibration data, it delivers the most stringent constraint on the AC model. 37Ar events with S2 areas smaller than 400 PE are selected to test the AC model. The predicted number of AC events was 731.6, while 733 events were observed in the data. The statistics in the AC model is very large so the uncertainty of the predicted AC rate is neglected in the following GOF tests. The 1D GOF tests in S1, S2, , , and the 2D test in S1–S2 all yield p-values between 0.05 and 0.95. The result of the 2D test with a p-value of 0.61 is shown in Fig. 8. The model was further validated with events removed by selections targeting the AC background inside the WIMP ROI. Similar tests were also performed on 220Rn calibration data. All these tests show the model and data to be compatible. Due to an over-fitting issue found in the GBDT training process, we conservatively estimated the rate uncertainty to be 30% for both the near- and far-wire regions. The expectation value of AC events in the WIMP search dataset in the ROI is .
@@ III.6 Surface Background Model
In XENONnT, the TPC wall is made of Polytetrafluoroethylene (PTFE), which is known to accumulate isotopes from the decay chain of atmospheric 222Rn, which decays down to 210Pb, an isotope with a half-life of 22.2 years. Radioactive decays on the wall surface can result in events with reduced S2 signals due to charge losses, which gives rise to a background that can leak into the WIMP signal region. Three components contribute to the surface background in the WIMP ROI: the two -decays of 210Pb and 210Bi, and the recoiling 206Pb following the -decay of 210Po. Independent screening measurements indicate a 210Po -decay activity on the PTFE surface of [54]. Due to the complexity of electric field conditions near the surface and the loss of ionization electrons to the detector walls, we employ a data-driven approach to model this background component in the space of cS1–cS2––.
The surface background model in the space of – was developed using 210Po -events. These events are mono-energetic (5.4 MeV) and thus easily identifiable through their characteristic S1 signal, yet they present a wide range of S2 sizes due to variable charge loss to the walls. The and distributions of 210Po -events were seen to match those of the lower-energy -events. The radial profiles were modeled using a Student’s generalized skew-t distribution. The radial distribution was fitted independently for different S2 sizes as shown in Fig. 9 to account for the S2-size dependent position resolution [11]. The profile was modeled using a linear function, to account for lower rates of surface events observed at the bottom of the TPC, likely due to the increased charge insensitivity near the walls at the bottom of the detector [7].
The cS1–cS2 distribution was modeled using a 2D Gaussian adaptive kernel density estimation (aKDE), built using events reconstructed outside the detector walls. The resulting model was then validated against the events reconstructed between the walls and the fiducial volume, in order to rule out any radial bias in the cS1–cS2 model. Figure 4 (lower left) shows the projection of the four-dimensional model on cS1–cS2 in the ROI. The absolute rate of events in the blinded region was inferred from the radial distributions of adjacent, non-blinded regions in cS1–cS2.
@@ IV.2 Nuisance Parameters and Constraints
| Relative signal efficiency | NR model uncertainty | ||
@@
In addition to our parameter of interest , we parameterize the space of background hypotheses using nuisance parameters in the likelihood function of Eq. 5. These parameters control both background rates and shape. The relative signal efficiency is a special parameter that quantifies the rate uncertainty of the NR emission model, as described in Sec. III.1. By construction, it takes the nominal value of 1.0 for an unconstrained fit, but it can depart from this value for fits with constrained cross sections, leading to a broadening of the confidence interval. In total, twelve nuisance parameters are defined, which are listed in Tab. 1. The uncertainties of the nominal expectation values correspond to the width of a Gaussian constraint term. The ancillary likelihood function is the product of all constraint terms. Most parameters are constrained via ancillary measurements, which are obtained from sidebands (e.g. data outside the ROI) or external measurements in combination with simulations. More details on the constraints on neutron, CENS, AC, and surface background rates were discussed in Sec. III. The ER rate is a free parameter in the fit and is fully constrained by the WIMP search data. The ER shape parameters obtained with PCA are constrained via the simultaneous fit of the calibration dataset via the term in Eq. 5.
@@ IV.3 Confidence Intervals and Discovery Significance
Knowing the distribution of under different hypotheses is essential for calculating discovery significances and confidence intervals. Due to the low rate of background resembling the signal, asymptotic formulae as discussed in [66] are not necessarily a good approximation. Therefore, we primarily use toy-MCs to estimate the distribution of the test statistic given that is the true cross section. We use the custom-developed blueice [67] framework to define the likelihood function. This Python package provides an efficient interpolation (“template morphing”) between PDFs evaluated for different discrete values of shape nuisance parameters. This allows shape parameters to be considered for which we have no analytical description.
For each signal model considered, we compute the discovery significance and the confidence interval by comparing the measured test statistic with the distribution under each hypothesis [68, 69]. Testing yields the (local) discovery significance, while confidence intervals are computed by finding the region where is rejected at a confidence level given the data, which is illustrated for three background-only toy-MCs (three black parabola-like lines) in Fig. 10 (top). The minimum of each curve corresponds to the respective best-fit value . Upper limits (ULs) and lower limits are obtained by finding the cross sections at which reaches the critical region. The threshold of the 90 CL critical region is precomputed as the 90th percentile of the test statistic distribution from toy-MCs with injected signal corresponding to the respective cross section. For low cross sections, it deviates from the asymptotic two-sided threshold indicated as a horizontal gray dotted line.
@@
Figure 11 shows the three-sigma discovery potential – the expected frequency of achieving a discovery with at least three-sigma significance – as a function of WIMP mass and cross section. The asymptotic threshold for a three-sigma discovery was applied. Contours indicating equal signal event counts in the nominal model reveal that the median three-sigma discovery potential corresponds to approximately 10 events at high WIMP masses, decreasing to about 6 events at lower masses.
Statistical fluctuations as well as mismodeling, such as overestimated background rates, can yield arbitrarily low ULs, which may result in the spurious exclusion of models beyond the experiment’s sensitivity. To mitigate this issue, various methods have been proposed [70, 71, 72]. The XENONnT SR0 WIMP search follows the recommendation of [48] to use power-constrained limits (PCLs) [72]. In this method, a signal size threshold is selected, at which the experiment has a “rejection power” – the probability of excluding a given signal under the background-only hypothesis. If an UL falls below this threshold, the threshold value is reported instead. These thresholds correspond to the quantiles of the UL distribution used to compute the sensitivity band illustrated in Fig. 10. For instance, choosing sets the threshold to the -1 sigma quantile of the band, while truncates ULs at the band’s median. In [48], the fiducial choice of was suggested. However, the choice was erroneously based on the discovery power, which corresponds to the probability of rejecting the background-only hypothesis given an alternative hypothesis with a specific signal size. In the absence of an updated recommendation, we chose the very conservative choice of for the first WIMP search of XENONnT.
@@ IV.4 Goodness of Fit and Blinding Procedure
To verify that the best-fit models describe our WIMP search data well, we defined a GOF test before unblinding, and after studies to confirm the power to reject various forms of mismodeling. The test uses a binned Poisson likelihood with 15 equiprobable bins in the cS1–cS2 space, defined from the model being tested. To compute a p-value, the distribution of the test statistic under the best-fit hypothesis is derived through toy-MCs. Specifying the test and acceptance threshold (90% CL) before unblinding ensured that the results were not influenced by statistical fluctuations expected from the low-statistic WIMP search dataset. For the background-only fit we found a p-value of and for a fit with an additional WIMP signal the p-value is . Both indicate no strong mismodeling in the predefined parameter space. Using GOF tests with well-studied power to discover signal-like mismodeling was also designed to replace the “safeguard” ER shape parameter defined in [73], due to its computational cost and susceptibility to some kinds of mismodeling.
When searching for new signals in data, it is crucial to avoid the experimenter’s bias on the result [74]. In our WIMP search, we adopt a common strategy for bias mitigation: blinding the signal region of the WIMP search dataset until all steps of the data analysis are finalized. Initially, events in the ER and NR bands of the WIMP search dataset with a reconstructed ER energy below were blinded. This involved defining the ER and NR bands in cS1 and cS2 based on quantiles in cS2 for slices in cS1 of calibration data. In the first step, all events above and those above the sigma quantile of the ER band were unblinded for the analysis presented in [65]. The signal region for the WIMP search (indicated in Fig. 5) remained blinded until the analysis procedure was finalized. The unblinded SR0 WIMP search data showed no significant excess for any of the tested WIMP masses with local discovery p-values . For this reason, we reported new ULs on the spin-independent WIMP-nucleon cross section across WIMP masses ranging from to , with a minimum of at .
@@ Acknowledgements.
We gratefully acknowledge support from the National Science Foundation, Swiss National Science Foundation, German Ministry for Education and Research, Max Planck Gesellschaft, Deutsche Forschungsgemeinschaft, Helmholtz Association, Dutch Research Council (NWO), Fundacao para a Ciencia e Tecnologia, Weizmann Institute of Science, Binational Science Foundation, Région des Pays de la Loire, Knut and Alice Wallenberg Foundation, Kavli Foundation, JSPS Kakenhi, JST FOREST Program, and ERAN in Japan, Tsinghua University Initiative Scientific Research Program, Domaine d’intérêt majeur en astrophysique et conditions d’apparition de la vie+ (DIM-ACAV+) Région Ile-de-France, and Istituto Nazionale di Fisica Nucleare. This project has received funding/support from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN. We gratefully acknowledge support for providing computing and data-processing resources of the Open Science Pool and the European Grid Initiative, in the following computing centers: the CNRS/IN2P3 (Lyon - France), the Dutch national e-infrastructure with the support of SURF Cooperative, the Nikhef Data-Processing Facility (Amsterdam - Netherlands), the INFN Centro Nazionale Analisi Fotogrammi (CNAF) (Bologna - Italy), the San Diego Supercomputer Center (San Diego - USA) and the Enrico Fermi Institute (Chicago - USA). We acknowledge the support of the Research Computing Center (RCC) at The University of Chicago for providing computing resources for data analysis. We are grateful to Laboratori Nazionali del Gran Sasso for hosting and supporting the XENON project.References
@@
- E. Aprile et al. [2024a] E. Aprile et al. (XENON collaboration), The XENONnT Dark Matter Experiment, Eur. Phys. J. C 84, 784 (2024a).
@@
- E. Aprile et al. [2024b] E. Aprile et al. (XENON collaboration), The neutron veto of the XENONnT experiment: Results with demineralized water, preprint (2024b), arXiv:2412.05264 [physics.ins-det] .
@@
- E. Aprile et al. [2025a] E. Aprile et al. (XENON collaboration), XENONnT analysis: Signal reconstruction, calibration and event selection, Phys. Rev. D 111, 062006 (2025a), arXiv:2409.08778 [hep-ex] .
@@
- E. Aprile et al. [2025b] E. Aprile et al. (XENON collaboration), Erratum: XENON1T Dark Matter Data Analysis: Signal & Background Models, and Statistical Inference [Phys. Rev. D 99, 112009 (2019)], Phys. Rev. D 111, 059904 (2025b).
@@
- Lenardo et al. [2019] B. G. Lenardo et al., Low-Energy Physics Reach of Xenon Detectors for Nuclear-Recoil-Based Dark Matter and Neutrino Experiments, Phys. Rev. Lett. 123, 231106 (2019).
@@
- Lindhard et al. [1963] J. Lindhard, M. Scharff, and H. E. Schioett, Range concepts and heavy ion ranges (notes on atomic collisions, II), Kgl. Danske Videnskab. Selskab. Mat. Fys. Medd. 33, 1 (1963).